OpenStudy (anonymous):

How fast must you drive to cover 340 miles in 6 hours? Round your answer to the nearest whole number. A. 55 mph B. 58 mph C. 60 mph D. 57 mph

OpenStudy (campbell_st):

speed = distance travalled /time taken

OpenStudy (anonymous):

D

OpenStudy (campbell_st):

nope... not D

OpenStudy (anonymous):

340/6=56.67=57, D

OpenStudy (campbell_st):

@vaibhav_raj one of the concepts of OpenStudy.. is to help... but not give answers

OpenStudy (anonymous):

From Mathematica:\[\text{Round}\left[\frac{340}{6},1\right]=57\text{ mph}\]